OpenStudy (anonymous):

Solve the following equation.Be sure to show all work. You should have a minimum of 7 lines of work. 5(2x - 7) = 33 - 3(6x + 4) Next, type out in words the steps you took to solve the equation

OpenStudy (anonymous):

Expand thx-e bracket and solve for x 10x-35=33-18x-12 10x+18x=33+35-12 28x=56 divide both side by 28 therfore x=2 hope that help.. :)